Bob is 20 years older than Ben. In 10 years, Bob will be twice as old as Ben. How old is each now?: Bob is 20 years older than Ben. In 10 years, Bob will be twice as old as Ben. How old is each now?
Answer by nerdybill(1051)

Bob is 20 years older than Ben. In 10 years, Bob will be twice as old as Ben. How old is each now?
.
Let x = age of Ben
then because "Bob is 20 years older than Ben"
x+20 = age of Bob
.
Our equation is derived from: "In 10 years, Bob will be twice as old as Ben."
(x+20) + 10 = 2(x+10)
x + 30 = 2x + 20
30 = x + 20
10 years = x (age of Ben)
.
x+20 = 10 + 20 = 30 years (age of Bob)